OpenStudy (anonymous):

Slope equation: passes thru (3,-1) perpendicular to y=4x+1 I can't reply back I'm on safari, but I will medal if your information helps me found out the answer!(:

zepdrix (zepdrix):

lol safari c: out hunting those lions and tigers and bears huh? \(\Large\rm y=\color{royalblue}{4}x+1\)

zepdrix (zepdrix):

This 4 is the slope of the given line. If the line we're searching for is `perpendicular` to this line, it will have a slope which is the `negative reciprocal` of this slope.

zepdrix (zepdrix):

So our m will be, \(\Large\rm m=-\frac{1}{4}\)

zepdrix (zepdrix):

We can write our line in the form, \(\Large\rm y=mx+b\) So we curently have, \(\Large\rm y=-\frac{1}{4}x+b\)

zepdrix (zepdrix):

Your line passes through the point (3,-1). So plug that point into our equation that we have so far, and solve for b. After you have found your b, go back to this, \(\Large\rm y=-\frac{1}{4}x+b\) and plug in the b. That will be your answer.

zepdrix (zepdrix):

That is the slope-intercept form of a line.